Company XYZ's CFO has provided the following information: The company's capital budget is expected to be $5,000,000. The company's target capital structure is 70 percent debt and 30 percent equity. The company's net income is $4,500,000. If the company follows a residual dividend policy, what portion of its net income should it pay out as dividends this year?
Answer #1

Amount of budget to be financed with equity = budget amount*equity% = 5000000*0.3=1500000

Amount paid as dividend = Net income-Amount of budget to be financed with equity = 4500000-1500000=

=3000000

%age paid out as dividend = dividend/net income = 3000000/4500000=66.67%